### v1

# Are you doing better than random guessing? A call for using negative controls when evaluating causal discovery algorithms

Abstract
New proposals for causal discovery algorithms are typically evaluated using simulations and a few select real data examples with known data generating mechanisms. However, there does not exist a general guideline for how such evaluation studies should be designed, and therefore, comparing results across different studies can be difficult. In this article, we propose a common evaluation baseline by posing the question: Are we doing better than random guessing? For the task of graph skeleton estimation, we derive exact distributional results under random guessing for the expected behavior of a range of typical causal discovery evaluation metrics (including precision and recall). We show that these metrics can achieve very large values under random guessing in certain scenarios, and hence warn against using them without also reporting negative control results, i.e., performance under random guessing. We also propose an exact test of overall skeleton fit, and showcase its use on a real data application. Finally, we propose a general pipeline for using random controls beyond the skeleton estimation task, and apply it both in a simulated example and a real data application.
1 Introduction
Causal discovery algorithms seek to infer information about a causal data generating mechanism by analyzing empirical data it generated. The causal data generating mechanism is typically represented by a causal graph, for example an equivalence class of directed acyclic graphs (DAGs). A highly productive research community has published a plethora of new causal discovery algorithms within the last 30 years or so. Naturally, this fast growing battery of available algorithms requires some standards and guidelines for evaluating and benchmarking their performance. Because the result of a causal discovery algorithm is an estimated graph (or family of graphs), rather than one or more scalars, it is not entirely straight-forward to use classic approaches for performance evaluation from neither machine learning nor statistics.
Nonetheless, machine learning classification metrics originally developed for evaluating prediction tasks, are often used to evaluate causal discovery algorithms. Most commonly, precision and recall, or possibly their harmonic mean, the F1 score, are reported, although some studies also focus on other metrics, e.g. negative predictive value (Petersen et al., 2023b). These metrics are computed from graph-level confusion matrices summarizing either agreement on placement of oriented edges (primarily used for DAG discovery evaluation), adjacencies (i.e., edge placement without considering orientation), and/or arrowheads among correctly placed adjacencies (conditional orientation). Typically, they are reported as averages over numerous simulations. Sometimes the results are stratified by graphical parameters (e.g., true graph density), data-related parameters (e.g., sample size), or simply reported as averages across several such settings.
Alternative metrics developed specifically for graphs also exist, where the structural hamming distance is the most widely used example, probably due to its cheap computation and easy interpretation. An alternative metric focusing more on the causal implications of the graphs is the structural intervention distance (SID) (Peters and Bühlmann, 2015), although it is most naturally suited for DAG-DAG comparison, and hence not readily applicable for all discovery evaluation tasks. A more recent proposal is the adjustment identification distance (Henckel et al., 2024), which also focuses on differences in causal inference based on the graph.
There are thus many different possible choices of metrics for evaluating causal discovery algorithms. However, no general guidelines exist for how to then interpret the values these metrics take. What is a high or low number? Of course, in a given evaluation study, multiple algorithms can be compared according to their relative performance. But it is not straight-forward to assess what values of the metrics should then be expected on just slightly different evaluation settings. This makes it very difficult to compare results across different evaluation studies with just marginally different designs.
In other sciences, e.g. epidemiology, negative controls are used to ensure the validity of a given study design (Lipsitch et al., 2010). A negative control is an experimental setting where the experimenters expect to see no effect. For example, we expect smoking to affect death by lung cancer, but probably not death by traffic accidents, and hence if we see an effect of smoking on death by traffic accidents in our study, there might be issues with the design. By analyzing data from the negative control scenario, and comparing it with the intended experiment, it may thus be possible to quantify unaccounted biases that are confounded by the experimental design. It is hence a pragmatic approach that acknowledges that we cannot always understand all aspects of a given scientific study design, and what may bias it, but we can instead sometimes specify what should not be able to affect our outcome of interest.
We propose to use a negative control concept to obtain an interpretable benchmark for any causal discovery evaluation study: Namely, to investigate what values of the metrics of interest can be obtained using random guessing (a negative control). We discuss random controls in two different settings: First, we consider the task of skeleton estimation, i.e., estimating e.g. a DAG without any orientation information. In this case, we derive exact distributional results for the expected behavior under random guessing (Section 3), and we use these results to compute expected negative control values for a range of often-used metrics (Section 4). We furthermore propose an exact test of overall skeleton fit (Section 5, and provide an example of its use of real data. Secondly, we consider more general metrics, that are not just concerned with skeleton estimation, and propose a negative control pipeline for this case (Section 6). We provide two examples of its use, both in a simulation study and in a real data application.
But before we turn to these general results, we provide an example of a case where well-known metrics such as adjacency precision and recall do perhaps not behave exactly as one would expect.
2 Precision and recall: A cautionary tale
a True DAG
[…]
b Estimated DAG
[…]
Figure 1: The true underlying DAG (a) and an estimated DAG (b) obtained using an undisclosed causal discovery procedure.
Consider the two DAGs in Figure 1. The graph on the left (a) is the true DAG, and the graph on the right (b) is an estimated graph produced by a causal discovery procedure. We compute their adjacency confusion matrix in order to evaluate the performance of the discovery procedure:
Truth
Adjacency Non-adjacency
Estimate Adjacency $tp=6$ $fp=1$
Non-adjacency $fn=2$ $tn=1$
which results in
[…]
Are these numbers high or low? Although these values are not too far off from the performance of well-established causal discovery algorithms on simulated data (and much better than typical performance on real data), we will argue that they are indeed as low as can be for this specific discovery task, since the applied ”discovery algorithm” was simply random guessing. The DAG in Figure 1 (b) was drawn randomly among DAGs with 7 edges over 5 nodes (Erdös-Rényi type graph), and hence had absolutely no information about the true data generating mechanism.
The example provided here is not just the result of a conveniently chosen random seed. The example graphs were in fact chosen specifically to match the most typical performance when using a randomly drawn 5 node DAG with 7 edges to estimate the skeleton of a ”true” 5 node DAG with 8 edges. This can for example be demonstrated in a simulation study: We simulated 1000 random Erdös-Rényi type DAGs over 5 nodes each with 7 edges and used these ”random guesses” as estimates of the DAG in Figure 1 (a). This resulted in a median precision of 0.86 and a median recall of 0.75, i.e., numbers that exactly match the performance of the example just described.
Is it then a curious artefact for very dense graphs? Or ”small” graphs over e.g. 5 nodes? This is not the case. As we will show in the following section, the phenomenon does not depend on the number of nodes, and depending on the choice of metrics, can occur also in modestly dense graphs.
3 Distributional results for adjacency metrics under random guessing
Truth
Adjacency Non-adjacency Total
Estimate Adjacency $TP$ $FP$ $m_{\text{est}}$
Non-adjacency $FN$ $TN$
Total $m_{\text{true}}$ $m_{\text{max}}$
Table 1: A generically labelled adjacency confusion matrix. Entries marked with dashes are sums that will not be used for the derivations here.
Consider a DAG $G$ over $d$ nodes, and let $m_{\text{true}}$ denote the number of edges in $G$. Let $\hat{G}$ be another DAG over $d$ nodes used as an estimate of $G$, and let $m_{\text{est}}$ be the number of edges in $\hat{G}$. Finally, let $m_{\text{max}}=\sum_{i=1}^{d-1}i=\frac{1}{2}(d-1)d$ denote the maximal number of possible edges in a DAG over $d$ nodes (corresponding to a fully connected graph).
We can describe the performance of $\hat{G}$ as an adjacency/skeleton estimator of $G$ through a (generic) confusion matrix as seen in Table 1. Note that for a causal discovery problem, $m_{\text{max}}$ and $m_{\text{true}}$ will generally be given a priori, as the number of nodes does not change, and the ground truth is considered fixed. Moreover, for many causal discovery procedures, it further makes sense to consider $m_{\text{est}}$ fixed – at least for a specific value of a tuning parameter (e.g. significance level for testing or penalty for a score) – as we do not often try to estimate the correct number of edges from data (see Sections 6 and 7 for considerations in cases where this latter assumption is not meaningful).
We now make the following important observation: If edges are placed uniformly in both $G$ and $\hat{G}$ (corresponding to an Erdös-Rényi type graph), edge placement can be seen as a random urn experiment with two colors of balls (say, blue and white), where white balls correspond adjacencies included in $G$, and blue balls are adjacencies not in $G$. A random causal discovery procedure will then metaphorically draw ”balls” (i.e., edges) randomly without replacement, and some will be correct (white), while others will be false (blue). Since the number of white balls ($m_{\text{true}}$), the number of draws ($m_{\text{est}}$) and the total number of balls ($m_{\text{max}}$) are all known a priori, the number of drawn white balls (true positive adjacencies) will by definition follow a hypergeometric distribution parameterized by $m_{\text{max}}$, $m_{\text{true}}$ and $m_{\text{est}}$:
$\text{TP}\operatorname{|}m_{\text{max}},m_{\text{true}},m_{\text{est}}\sim\text{HyperGeom}(m_{\text{max}},m_{\text{true}},m_{\text{est}})$
Note that this is an exact distributional result, not an asymptotic statement.
This observation gives rise to several useful applications: First, we can compute the expected value, median and uncertainty estimates (e.g. confidence interval) for the number of true positive adjacencies under random guessing. Secondly, since we are also conditioning on $m_{\text{max}}$, $m_{\text{true}}$ and $m_{\text{est}}$, we can further compute expectations and draw statistical inference for any function of the confusion matrix, including precision, recall and F1. We provide formulas for these below in Section 4. Thirdly, we can construct an exact statistical test of overall skeleton fit by considering how much the number of true positives in a given estimated graph diverts from its expected distribution under a null hypothesis of random edge placement. We propose such a test in Section 5 below.
4 Expectations and quantiles of adjacency metrics under random guessing
Metric Expected value Quantile
Precision $\frac{m_{\text{true}}}{m_{\text{max}}}$ $\frac{k_{q}}{m_{\text{est}}}$
Recall $\frac{m_{\text{est}}}{m_{\text{max}}}$ $\frac{k_{q}}{m_{\text{true}}}$
F1 $\frac{2\cdot m_{\text{est}}\cdot m_{\text{true}}}{m_{\text{max}}\cdot m_{\text{est}}+m_{\text{max}}\cdot m_{\text{true}}}$ $\frac{2\cdot k_{q}}{m_{\text{est}}+m_{\text{true}}}$
NPV $1-\frac{m_{\text{true}}}{m_{\text{max}}}$ $\frac{m_{\text{max}}-m_{\text{est}}-m_{\text{true}}+k_{q}}{m_{\text{max}}-m_{\text{est}}}$
Specificity $1-\frac{m_{\text{est}}}{m_{\text{max}}}$ $\frac{m_{\text{max}}-m_{\text{est}}-m_{\text{true}}+k_{q}}{m_{\text{max}}-m_{\text{true}}}$
Table 2: Expected values and quantile expressions under random guessing for five commonly used metrics computed for adjacency performance. $q_{k}$ denotes the $k$th quantile from $\text{HyperGeom}(m_{\text{max}},m_{\text{true}},m_{\text{est}})$.
Since $TP\operatorname{|}m_{\text{max}},m_{\text{true}},m_{\text{est}}\sim\text{HyperGeom}(m_{\text{max}},m_{\text{true}},m_{\text{est}})$, by definition we have that
$\text{E}(TP\operatorname{|}m_{\text{max}},m_{\text{true}},m_{\text{est}})=\frac{m_{\text{est}}\cdot m_{\text{true}}}{m_{\text{max}}}$
and by considering the quantile function of $\text{HyperGeom}(m_{\text{max}},m_{\text{true}},m_{\text{est}})$, we can find the expected median, or construct a confidence interval as e.g. the central 95% of the distribution.
Table 2 provides an overview of expected values and quantiles under random guessing for five metrics commonly used for evaluating adjacency placement for causal discovery algorithms, namely precision, recall, F1 score, negative predictive value (NPV) and specificity. As an example, we here showcase derivations for precision, and refer to Appendix A for derivations for the remaining four metrics.
Expectation and quantiles for adjacency precision
We first express precision as a function of $TP,m_{\text{max}},m_{\text{true}},m_{\text{est}}$:
$\text{precision}\operatorname{|}(m_{\text{max}},m_{\text{true}},m_{\text{est}})$ $=\frac{TP}{TP+FP}$ (1)
$=\frac{TP}{TP+m_{\text{est}}-TP}$ (2)
$=\frac{TP}{m_{\text{est}}}$ (3)
Since this is a linear function of $TP$, we can straight-forwardly compute the expectation:
$\text{E}(\text{precision}\operatorname{|}m_{\text{max}},m_{\text{true}},m_{\text{est}})$ $=\frac{1}{m_{\text{est}}}\text{E}(TP\operatorname{|}m_{\text{max}},m_{\text{true}},m_{\text{est}})$
$=\frac{m_{\text{true}}}{m_{\text{max}}}$
The linearity in equation (3) also makes it easy to obtain e.g. a 95% exact confidence interval under the null hypothesis of random guessing, by simply applying the transformation from (3) to the appropriate quantiles of $\text{HyperGeom}(m_{\text{max}},m_{\text{true}},m_{\text{est}})$. For example, an exact 95% confidence interval for precision under the null is given by
$\text{CI}_{0.95}=\left(\frac{1}{m_{\text{est}}}q_{(0.025,m_{\text{max}},m_{\text{true}},m_{\text{est}})},\frac{1}{m_{\text{est}}}q_{(0.975,m_{\text{max}},m_{\text{true}},m_{\text{est}})}\right)$
where $q_{(k,m_{\text{max}},m_{\text{true}},m_{\text{est}})}$ is used to denote the $k$th quantile of the probability mass function of $\text{HyperGeom}(m_{\text{max}},m_{\text{true}},m_{\text{est}})$. Similarly, we obtain the median precision by simply computing
$\text{median}(\text{precision}\operatorname{|}m_{\text{max}},m_{\text{true}},m_{\text{est}})=\frac{1}{m_{\text{est}}}q_{(0.5,m_{\text{max}},m_{\text{true}},m_{\text{est}})}$
A notable feature of Table 2 is that, conditional on $(m_{\text{max}},m_{\text{true}},m_{\text{est}})$, the expected precision is simply the density of the true DAG $G$, and the expected recall is the density of the estimated DAG $\hat{G}$. Furthermore, the expected values of NPV and specificity are given as 1 minus the expectations of precision and recall, respectively, and hence they do not provide additional information. However, without random guessing, this is of course not generally the case, so they are still useful to compute in order to provide a nuanced and multifaceted evaluation of a given causal discovery procedure.
Moreover, we note that under random guessing, the expected precision does not depend on the number of edges in the estimated graph ($m_{\text{est}}$), only on the number of edges in the true graph ($m_{\text{true}}$) and the maximal possible number of edges ($m_{\text{max}}$). But recall increases linearly as a function of the number of estimated edges. Hence, if we are using random guessing, a ”free lunch” in optimizing precision and recall can be had by simply estimating a very large number of edges, even including the trivial fully connected graph. This can also be seen from the expected value of the F1 score under random guessing, which increases monotonically with the number of estimated edges: It is always better to just add another edge.
We will now consider two small example applications of the results from Table 2. First, we revisit the example from Section 2 and compute the median, expected value, and a 95% confidence interval for precision and recall for this case. Next, we provide an overview of how the expected F1 score varies as a function of $m_{\text{est}}$ and $m_{\text{true}}$ under random guessing.
4.1 Example: Expected precision and recall for a dense 5 node DAG skeleton
Consider the problem from Section 2 regarding estimating a DAG skeleton over 5 nodes. Such a DAG can have at most $m_{\text{max}}=\frac{1}{2}(5-1)5=10$ edges. Assume that the true DAG has $m_{\text{true}}=8$ edges, while a randomly drawn graph over the same 5 nodes has $m_{\text{est}}=7$ edges. What performance can we then expect from this random guessing procedure? With reference to Table 2, we find
$\text{E}(\text{precision}\operatorname{|}(m_{\text{max}}=10,m_{\text{true}}=8,m_{\text{est}}=7))=\frac{8}{10}=0.80$
with a 95% confidence interval of
$\left(\frac{q_{(0.025,10,8,7)}}{7},\frac{q_{(0.975,10,8,7)}}{7}\right)=\left(\frac{5}{7},\frac{7}{7}\right)=(0.71,1.00).$
Hence for this DAG estimation task, it will not be highly unusual to obtain adjacency precisions as high as 1.00 under random guessing, and hence adjacency precision is not very useful for assessing performance. We can also compute the median:
$\text{median}(\text{precision})=\frac{q_{(0.5,10,8,7)}}{7}=\frac{6}{7}=0.86$
This is the same value as found in the simulations presented in Section 2.
Similarly, for adjacency recall we find
$\text{E}(\text{recall}\operatorname{|}(m_{\text{max}}=10,m_{\text{true}}=8,m_{\text{est}}=7))=\frac{7}{10}=0.70$
and we compute a 95% confidence interval as
$\left(\frac{k_{0.025,10,8,7}}{8},\frac{k_{0.975,10,8,7}}{8}\right)=\left(\frac{5}{8},\frac{7}{8}\right)=(0.63,0.88).$
One is not included in this confidence interval and hence adjacency recall does have some discriminatory power for this DAG estimation task. We compute the median:
$\text{median}(\text{recall})=\frac{q_{(0.5,10,8,7)}}{8}=\frac{6}{8}=0.75$
Once again, this matches our simulation-based findings from Section 2.
4.2 Example: F1 scores for a 5 node DAG with varying density
Figure 2: Expected F1 scores obtained under random guessing for estimating a 5 node DAG with varying true density ($m_{\text{true}}$, displayed on the horizontal axis) and varying density of the estimate ($m_{\text{est}}$, marked in color).
Figure 2 provides an overview of obtained F1 scores under random guessing across all possible combinations of estimated number of edges (marked in color) and true number of edges (horizontal axis) for 5 node DAGs. We see that it is quite possible to obtain a large F1 by random guessing if the true DAG is not very sparse, and especially, if the estimate also is not very sparse. But we stress that neither has to be overly or unrealistically dense either: For a true graph that has just 5 edges – i.e. the sparsest graph that is connected – a randomly drawn DAG with 5 edges will result in an expected F1 of 0.5, and placing all 10 possible edges results in an F1 score of 0.66. If we instead consider a more dense graph, e.g. a true DAG with 8 edges, we are back in then scenario already considered above, and we see that we can find a peak F1 score of 0.89 by placing all edges.
5 A test of overall skeleton fit
We can also use the distributional results presented above to construct an exact test of overall skeleton fit. More specifically, for an estimated DAG $\hat{G}$ with $m_{\text{est}}$ edges, we test the null hypothesis
$H_{0}:\hat{G}\text{ was obtained by randomly placing }m_{\text{est}}\text{ edges.}$.
This is done by comparing the observed number of true positives, $TP_{obs}$ with the appropriate hypergeometric distribution. Formally, if we let $X\sim\text{HyperGeom}(m_{\text{max}},m_{\text{true}},m_{\text{est}})$, a one-sided p-value for $H_{0}$ is computed as
$P(X\geq TP_{obs})$
i.e., the probability of getting as least as many true positives as the observed number, if edges were in fact randomly placed.
Note than since the text is exact (and based on a discrete probability distribution), it will be conservative.
5.1 Application on Metropolit cohort dataset
Experts
Adjacency Non-adjacency
TPC Adjacency 10 20
Non-adjacency 20 181
Table 3: Adjacency confusion matrix replicated from Petersen et al. (2023a).
We will reanalyze data from Petersen et al. (2023a). Here, the temporal PC algorithm (TPC) was used on a cohort data set of $n=3145$ Danish men to identify possible causes of depression and heart disease, as well as their interplay. Two experts were also asked to construct a model for the data based on existing studies and subject-field knowledge, and their DAG was compared to the output of TPC. For the comparison here, we assume that the expert model is correct and wish to evaluate if TPC performs better than a random control at estimating the expert model. The DAGs have 22 nodes, and hence $m_{\text{max}}=231$ possible edges.
Table 3 shows the adjacency confusion matrix comparing the expert and TPC models. Note that the two models did not disagree on edge orientation among shared adjacencies (although one shared adjacency was left unoriented by TPC). Hence in this case, the adjacency performance comparison summarizes all edge-wise comparisons of the two outputs. Note also that TPC was set to find the same number of edges as the experts did, i.e. $m_{\text{true}}=m_{\text{est}}=30$, and hence the symmetry in the confusion matrix is by design.
We conduct an overall test of skeleton fit by comparing the obtained number of true positives, $\text{TP}_{\text{obs}}=10$, with $\text{HyperGeom}(231,30,30)$ and we find $p=0.002$. Hence, we reject $H_{0}$ and conclude that TPC performs significantly better than random guessing in this application.
6 Simulation-based negative controls for more general metrics
Although the results provided above cover some of the most commonly reported metrics for causal discovery evaluation, other interesting metrics cannot be expressed as functions of the adjacency confusion matrix, and hence are out the scope of the results presented thus far.
One example is conditional orientation metrics (also sometimes referred to as arrowhead metrics) (see e.g. Andrews et al. (2019)). These metrics describe correct orientation among correctly placed edges. We conjecture that simple exact distributional results under random guessing do not exist for this classification task. The main issue is that placement of oriented edges does not correspond to an urn experiment analogue, as consecutive placement steps are not independent when the goal is to output e.g. a valid DAG: If we have already placed oriented edges such that $X\to Y\to Z$, it is (no longer) possible to have an edge pointing from $Z$ to $X$, as this would introduce a cycle and the graph would thus no longer be a valid DAG. Thus, describing expected behavior under random guessing when also taking edge orientations into account is more complicated.
However, we can easily use simulation to obtain an empirical estimate of the distribution of a given metric under random edge placement – oriented or not. We propose the following procedure:
1. Standard simulation study:
Conduct the simulation study as usual (focusing on one or more metrics of interest) and for each simulation repetition, store the true (simulated) DAG as well as the density (or number of edges) for the estimated DAG.
2. Negative control estimation:
Draw a large (e.g. 1000) number of random DAGs with densities sampled from the distribution obtained in Step 1.
3. Negative control evaluation:
Compare each of the random DAGs to the true DAGs from step 1 by computing the metric(s) of interest. Report the mean(s) as the expected performance under random guessing, and use the empirical quantiles to construct e.g. a 95% confidence interval.
4. Comparison:
Finally, compare the metrics obtained under random guessing with what is obtained from the algorithm being evaluated. In order to draw statistical inference, compute their pairwise performance differences for each ”true” DAG, and construct e.g. a 95% confidence interval using the empirical quantiles of the distribution of these differences.
We here focus on the case of a simulation study, where many different ground truths are simulated, but in Section 6.2 we also provide an example of how to adapt the procedure to be suited for evaluation of a real data application where there is only a single ground truth.
Note that it is important for obtaining valid statistical inference that it is conducted on the pair-wise differences in performance, as inference e.g. based on whether or not confidence intervals overlap is highly conservative (Knol et al., 2011).
6.1 Example: Structural Hamming distance in simulation study
The structural Hamming (SHD) distance measures the number of edge reversals, removals and additions in order to transform one graph (estimate) into another (truth). This metric hence includes both adjacency information and orientation information, and thus the results from Section 3 do not apply. We will instead compute an SHD negative control for evaluating the PC algorithm by use of simulation. We choose a simulation scenario, where it is well-known that PC struggles, namely a dense true data generating mechanism, and hence we do not expect PC to do better than a random control.
We proceed as follows (following the steps from above):
1. Standard simulation study:
We draw 1000 random graphs over 10 nodes, each with 30 edges (i.e., quite dense). From each DAG, we simulate 400 independent Gaussian observations with randomly drawn regression parameters and error variances [^1] We use default options for regression parameters and error variances from the simGausFromDAG() function in the causalDisco R package (Petersen, 2022).. For each dataset, we apply the PC algorithm with test significance level $\alpha=0.05$.
  1. (a)
    The metric of interest, SHD, is computed for each of the 1000 graphs, and we find a mean SHD of $27.3$ with a 95% confidence interval of $(21;33)$ (based on the empirical quantiles).
  2. (b)
    We store the ”true” DAGs as well as the distribution of number of edges across the 1000 applications of the PC algorithm. The number of edges range from 7 to 19 with a mean of $13.4$.
2. Negative control estimation:
We draw 1000 random DAGs over 10 nodes with number of edges independently sampled from the distribution from Step 1 (b).
3. Negative control evaluation:
For each of the 1000 negative controls, we compare with the ”true” DAGs from Step 1 by computing the SHD. We find a mean SHD of $30.1$ (95% CI: $24;35$) for the negative controls.
4. Comparison:
We find the mean (pairwise) difference in SHD for PC vs. random controls to be $2.8$, 95% CI: $(-5;11)$. As 00 is included in this confidence interval, we conclude that PC is not performing statistically significantly different from a random control in this scenario (at a 5% test level).
6.2 Application: Negative control for evaluation of NOTEARS on Sachs data
In this example, we evaluate a causal discovery procedure using a (single) real data set with a known ground truth. Hence, we only have a single ”true” DAG (and a single value of $m_{\text{est}}$), but we will again use repeated negative control simulations in order to gain more stable results and allow for computing a confidence interval.
We consider evaluation of the NOTEARS algorithm on the Sachs dataset (Sachs et al., 2005), which is often used to evaluate causal discovery algorithms. The ground truth DAG for the Sachs dataset has 11 nodes and 20 edges (although there also exists an alternative version with only 17 edges) [^2] We use this ”truth” version of the Sachs dataset because that is also what was used to evaluate the NOTEARS algorithm. We obtain the ground truth from the repository https://github.com/cmu-phil/example-causal-datasets.
In the original article proposing NOTEARS, the authors report that NOTEARS obtains an SHD of 22 on the Sachs dataset by placing 16 edges (Zheng et al., 2018). To compute performance on a negative control for this task, we simulate 1000 random DAGs over 11 nodes with 16 edges (i.e., same number of edges as NOTEARS). The average SHD over the random controls, comparing to the Sachs ground truth, is $27.4$ with a 95% confidence interval of $(22,33)$. Hence, testing at a 5% level, NOTEARS does not obtain significantly better SHD performance than the random control on the Sachs dataset (although the value lies on the border of the confidence interval).
7 Discussion
The distributional results for adjacency metrics presented in Section 3 are conditional on three quantities: The maximal number of edges in the DAG ($m_{\text{max}}$), the number of edges in the true DAG ($m_{\text{true}}$) and the number of edges in the estimated DAG ($m_{\text{est}}$). Clearly, conditioning on the first two is completely uncontroversial: $m_{\text{max}}$ is a mathematical property of the graph, and in order to do causal discovery, we have to assume that there exists an (unknown) true DAG with a set number of edges ($m_{\text{true}}$). But conditioning on $m_{\text{est}}$ may be debated. Our motivation for doing so is as follows: Many causal discovery algorithms – e.g. PC (Spirtes and Glymour, 1991), FCI (Spirtes et al., 2000), GES (Chickering, 2002), and GRaSP (Lam et al., 2022) – require choosing a tuning parameter, which will in practice directly control the number of outputted edges (e.g. test significance level in constraint-based algorithms or score penalties in score-based algorithms). The way causal discovery algorithms are mostly applied, this tuning parameter is not chosen in a data-driven manner, but rather set at somewhat arbitrary ”standard” values. Alternatively, in some instances, it may be chosen based on external background knowledge (Petersen et al., 2023a). In either case, the number of edges in the estimated graph is de facto chosen a priori, in which case we can meaningfully condition on it.
Some work has been proposed for data-driven tuning of causal discovery algorithms (Biza et al., 2020). If such methods are applied, the density of the estimated graph will generally be estimated from data. In this case, we lose the distributional results for the adjacency metrics presented in Section 3, as the number of true positives will no longer be hypergeometrically distributed. But we can still use the proposed simulation-based pipeline from Section 6 to obtain a negative control for such algorithms as well.
Another reason for preferring to condition on $m_{\text{est}}$ is related to interpretability. By only considering one value of $m_{\text{est}}$, we compare our causal discovery procedure with a well-specified negative control experiment, namely placing the same number of edges at random. This argument also has implications for how we ought to compare two different causal discovery algorithms, namely, by tuning them to estimate the same number of edges and then compare their outputs. Otherwise, we may be comparing sparse outputs with dense ones, and as we have seen above, such a comparison may not be meaningful, and will definitely be difficult to interpret. Moreover, by tuning them to produce equally dense outputs, we also remove tuning parameters from the evaluation equation altogether, and replace this parameter with the simpler and more interpretable notion of outputted graph density.
However, we want to stress that it can be problematic to consider only a single density in a simulation-based evaluation study if the algorithm being evaluated is able to thereby learn the (unique) intended density (Petersen et al., 2023b). This evaluation design flaw has been present for several supervised discovery algorithms (Li et al., 2020; Xu and Xu, 2021; Yu et al., 2019), and could harm transportability greatly. We propose that a range of densities should hence always be considered when conducting simulation-based evaluation studies of discovery algorithms that may learn the density directly from training data. But to ease interpretation, the results should ideally be presented stratified according to density.
The work presented here has focused on Erdös-Rényi type graphs. This assumption is important for the distributional results in Sections 3-5, as the hypergeometric distribution requires random draws. Non-central versions of the hypergeometric distribution allows for biased draws of edges, but we do not believe this is very useful for describing causal graphs: It would allow certain edges to be more likely to be present than others, but would still not consider graph properties beyond singular edges and hence not be appropriate for describing e.g. graphs that exhibit clustering. However, if a specific evaluation study wants to target such graphs, the simulation-based method proposed in Section 6 can straightforwardly be applied, simply by simulating random graphs from the intended target graph type.
As mentioned in Section 5, the exact distributional results for adjacency metrics will by definition result in conservative statistical inference, i.e. conservative control of type I error in statistical tests and overly wide confidence intervals. Due to the discrete nature of the hypergeometric distribution, this is especially pronounced when $m_{\text{max}}$ is small, i.e. when there are only few nodes. However, we argue that the considered null hypothesis is very crude – assuming completely random replacement of $m_{\text{est}}$ edges – and hence we do not consider conservative inference to be very problematic. Informally, we would ideally like to perform markedly better than random guessing, not just borderline significantly so!
In conclusion, we believe the results and examples provided here showcase that we need to acknowledge that causal discovery is not just another machine learning problem. Estimating a high-dimensional object such as a graph is difficult, and evaluating how well one did is equally challenging. A graph is more than the sum of its edges, and if we do not take into account the most fundamental property of the graphs we simulate for evaluation – its density – we are not producing useful results that will be likely to generalize to new data with other densities. We believe that the use of negative controls will be a useful next step in the direction of more transparent and interpretable evaluations. We of course all hope to do better than random guessing, so let us make it easy to see when we do - and when we do not.
Appendix A: Computations for Table YY
Below, we let $q_{(k,m_{\text{max}},m_{\text{true}},m_{\text{est}})}$ be the $k$th quantile from $\text{HyperGeom}(m_{\text{max}},m_{\text{true}},m_{\text{est}})$. We also note, and use repeatedly below, that $TP\operatorname{|}(m_{\text{max}},m_{\text{true}},m_{\text{est}})\sim\text{HyperGeom}(m_{\text{max}},m_{\text{true}},m_{\text{est}})$, and hence $\text{E}(TP\operatorname{|}m_{\text{max}},m_{\text{true}},m_{\text{est}})=\frac{m_{\text{est}}\cdot m_{\text{true}}}{m_{\text{max}}}$.
Recall:
We write recall as a function of $TP,m_{\text{max}},m_{\text{true}}$ and $m_{\text{est}}$:
recall $=\frac{TP}{TP+FN}$
$=\frac{TP}{TP+m_{\text{true}}-TP}$
$=\frac{TP}{m_{\text{true}}}$
Since this is a linear function of $TP$, the conditional expectation given ($m_{\text{max}},m_{\text{true}}$, $m_{\text{est}}$) is
$\text{E}(\text{recall}\operatorname{|}m_{\text{max}},m_{\text{true}},m_{\text{est}}))$ $=\frac{1}{m_{\text{true}}}\text{E}(TP\operatorname{|}m_{\text{max}},m_{\text{true}},m_{\text{est}})$
$=\frac{1}{m_{\text{true}}}m_{\text{est}}\frac{m_{\text{true}}}{m_{\text{max}}}$
$=\frac{m_{\text{est}}}{m_{\text{max}}}$
and the $k$th quantile of the recall distribution, conditional on $(m_{\text{max}},m_{\text{true}},m_{\text{est}})$, is given by
$\frac{1}{m_{\text{true}}}q_{(k,m_{\text{max}},m_{\text{true}},m_{\text{est}})}$
F1:
We write the F1 score as a function of $TP,m_{\text{max}},m_{\text{true}}$ and $m_{\text{est}}$:
F1 $=\frac{2\cdot TP}{2\cdot TP+FP+FN}$
$=\frac{2\cdot TP}{2\cdot TP+m_{\text{est}}-TP+m_{\text{true}}-TP}$
$=\frac{2\cdot TP}{m_{\text{est}}+m_{\text{true}}}$
Since this is a linear function of $TP$, the conditional expectation given ($m_{\text{max}},m_{\text{true}}$, $m_{\text{est}}$) is
[…]
$=\frac{2\cdot\frac{m_{\text{est}}\cdot m_{\text{true}}}{m_{\text{max}}}}{m_{\text{est}}+m_{\text{true}}}$
$=\frac{2\cdot m_{\text{est}}\cdot m_{\text{true}}}{m_{\text{max}}\cdot(m_{\text{est}}+m_{\text{true}})}$
and the $k$th quantile of the F1 distribution, conditional on $(m_{\text{max}},m_{\text{true}},m_{\text{est}})$, is given by
$\frac{2\cdot q_{(k,m_{\text{max}},m_{\text{true}},m_{\text{est}})}}{m_{\text{est}}+m_{\text{true}}}$
NPV:
We write the negative predictive value (NPV) as a function of $TP,m_{\text{max}},m_{\text{true}}$ and $m_{\text{est}}$:
NPV $=\frac{TN}{TN+FN}$
$=\frac{m_{\text{max}}-m_{\text{est}}-m_{\text{true}}+TP}{m_{\text{max}}-m_{\text{est}}-m_{\text{true}}+TP+FN}$
$=\frac{m_{\text{max}}-m_{\text{est}}-m_{\text{true}}+TP}{m_{\text{max}}-m_{\text{est}}}$
Since this is a linear function of $TP$, the conditional expectation given ($m_{\text{max}},m_{\text{true}}$, $m_{\text{est}}$) is
$\text{E}(\text{NPV}\operatorname{|}m_{\text{max}},m_{\text{true}},m_{\text{est}}))$ $=\frac{m_{\text{max}}-m_{\text{est}}-m_{\text{true}}+\text{E}(TP\operatorname{|}m_{\text{max}},m_{\text{true}},m_{\text{est}})}{m_{\text{max}}-m_{\text{est}}}$
$=\frac{m_{\text{max}}-m_{\text{est}}-m_{\text{true}}+\frac{m_{\text{est}}\cdot m_{\text{true}}}{m_{\text{max}}}}{m_{\text{max}}-m_{\text{est}}}$
$=1-\frac{m_{\text{true}}}{m_{\text{max}}}$
$=1-\text{E}(\text{precision}\operatorname{|}m_{\text{max}},m_{\text{true}},m_{\text{est}})$
and the $k$th quantile of the NPV distribution, conditional on $(m_{\text{max}},m_{\text{true}},m_{\text{est}})$, is given by
$\frac{m_{\text{max}}-m_{\text{est}}-m_{\text{true}}+q_{(k,m_{\text{max}},m_{\text{true}},m_{\text{est}})}}{m_{\text{max}}-m_{\text{est}}}$
Specificity:
We write specificity as a function of $TP,m_{\text{max}},m_{\text{true}}$ and $m_{\text{est}}$:
specificity $=\frac{TN}{TN+FP}$
$=\frac{m_{\text{max}}-m_{\text{est}}-m_{\text{true}}+TP}{m_{\text{max}}-m_{\text{est}}-m_{\text{true}}+TP+FP}$
$=\frac{m_{\text{max}}-m_{\text{est}}-m_{\text{true}}+TP}{m_{\text{max}}-m_{\text{true}}}$
Since this is a linear function of $TP$, the conditional expectation given ($m_{\text{max}},m_{\text{true}}$, $m_{\text{est}}$) is
$\text{E}(\text{specificity}\operatorname{|}m_{\text{max}},m_{\text{true}},m_{\text{est}}))$ $=\frac{m_{\text{max}}-m_{\text{est}}-m_{\text{true}}+\text{E}(TP\operatorname{|}m_{\text{max}},m_{\text{true}},m_{\text{est}})}{m_{\text{max}}-m_{\text{true}}}$
$=\frac{m_{\text{max}}-m_{\text{est}}-m_{\text{true}}+\frac{m_{\text{est}}\cdot m_{\text{true}}}{m_{\text{max}}}}{m_{\text{max}}-m_{\text{true}}}$
$=1-\frac{m_{\text{est}}}{m_{\text{max}}}$
$=1-\text{E}(\text{recall}\operatorname{|}m_{\text{max}},m_{\text{true}},m_{\text{est}})$
and the $k$th quantile of the specificity distribution, conditional on $(m_{\text{max}},m_{\text{true}},m_{\text{est}})$, is given by
$\frac{m_{\text{max}}-m_{\text{est}}-m_{\text{true}}+q_{(k,m_{\text{max}},m_{\text{true}},m_{\text{est}})}}{m_{\text{max}}-m_{\text{true}}}$